a bag 10 white, 15 red and 8 green balls. A single draw of 3 balls is made. what is the probability that a white, a red and green balls are drawn


1
Expert's answer
2021-12-14T16:20:46-0500
"10+15+8=33"


"P(1W\&1B\&1G)=\dfrac{\dbinom{10}{1}\dbinom{15}{1}\dbinom{8}{1}}{\dbinom{33}{3}}=\dfrac{10(15)(8)}{5456}"

"=\dfrac{75}{341}\approx0.22"